OpenStudy (anonymous):

Angle DEF and angle FEG are supplementary. The measure of angle DEF = (9x +1) and the measure of angle FEG = (8x + 9). What is the measure of angle FEG?

Directrix (directrix):

By definition, the measures of supplementary angles sum to 180. m<DEF + m<FEG = 180 9x + 1 + 8x + 9 =180 17x + 10 = 180 17x = 170 x = 10 measure of angle FEG = (8x + 9) = 8*10 + 9 = 80 + 9 = 89 -->*